# with Self-Reflective Rationales

### v1

@@ Abstract
Large language models (LLMs) often generate inaccurate or fabricated information and generally fail to indicate their confidence, which limits their broader applications. Previous work elicits confidence from LLMs by direct or self-consistency prompting, or constructing specific datasets for supervised finetuning. The prompting-based approaches have inferior performance, and the training-based approaches are limited to binary or inaccurate group-level confidence estimates. In this work, we present the advanced SaySelf, a training framework that teaches LLMs to express more accurate fine-grained confidence estimates. In addition, beyond the confidence scores, SaySelf initiates the process of directing LLMs to produce self-reflective rationales that clearly identify gaps in their parametric knowledge and explain their uncertainty. This is achieved by using an LLM to automatically summarize the uncertainties in specific knowledge via natural language. The summarization is based on the analysis of the inconsistency in multiple sampled reasoning chains, and the resulting data is utilized for supervised fine-tuning. Moreover, we utilize reinforcement learning with a meticulously crafted reward function to calibrate the confidence estimates, motivating LLMs to deliver accurate, high-confidence predictions and to penalize overconfidence in erroneous outputs. Experimental results in both in-distribution and out-of-distribution datasets demonstrate the effectiveness of SaySelf in reducing the confidence calibration error and maintaining the task performance. We show that the generated self-reflective rationales are reasonable and can further contribute to the calibration. The code is made public at https://github.com/xu1868/SaySelf.
@@ 1 Introduction
While large language models (LLMs) exhibit remarkable proficiency in reasoning and generating effective responses [28, 37, 17, 39], they often produce fabricated information (a.k.a, hallucination) and typically hesitate to indicate their uncertainty when faced with unfamiliar questions [45, 23]. Determining how to accurately obtain reliable confidence estimates from LLMs is essential [42, 51], particularly when the responses are not limited to single tokens [^1] The confidence of a single-token answer can be derived from its token probability, which typically exhibits high calibration [28, 14, 7]..
@@
Previous work on eliciting confidence from LLMs includes prompting-based and training-based approaches. Prompting-based methods employ specific prompts to generate confidence scores or use answer consistency as a confidence indicator, though these can have poor calibration performance or significantly increase inference latency [36, 42]. Training-based approaches develop specialized datasets for fine-tuning that encourage LLMs to express confidence. However, these methods often provide suboptimal or binary confidence estimates, failing to accurately reflect the models’ confidence levels [22, 47].
@@
In this work, we present SaySelf, a training framework that teaches LLMs to generate more accurate and fine-grained confidence estimates. Importantly, SaySelf goes beyond the confidence elicitation in previous work, and further enables LLMs to generate self-reflective rationales that indicate their knowledge gap and explain their confidence estimates (Figure 1). We accomplish this by automatically generating a model-specific dataset for supervised fine-tuning using an off-the-shelf LLM (e.g., GPT-4 [28]). Specifically, for each question, we sample multiple reasoning chains from LLMs. We then perform clustering of the reasoning chains based on the semantic similarity and retain one instance per cluster. We instruct GPT-4 to analyze the selected instances from various clusters, and summarize the uncertainties regarding specific knowledge in natural language from a first-person perspective.
@@
For accurate confidence estimates, we employ reinforcement learning to calibrate LLMs’ confidence estimate in each response. We design a reward function that incentivizes LLMs to produce accurate, high-confidence predictions and imposes penalties for overconfidence in incorrect responses.
@@
In experiments, we evaluate SaySelf on multiple knowledge-extensive question-answering tasks. We show that SaySelf significantly reduces the confidence calibration error and maintains the task performance. In addition, the generated self-reflective rationales effectively capture the internal uncertainty and can further improve the calibration performance.
@@ 2 Related Work
LLMs’ hallucination refers to instances where these models generate information that is not supported by their training data or the input provided [48, 20, 1]. Numerous research is dedicated to exploring the causes of hallucination [9, 27, 13] and developing methods to detect or mitigate hallucination [38, 31, 3]. Besides the hallucination, the reluctance of LLMs to express uncertainty when they are unable to solve tasks can further erode trust in these systems [16, 50]. Existing research identifies the tendency in LLMs to fabricate information when addressing unknown questions [23, 15, 2]. This inability can be traced back to the supervised instruction finetuning (SFT) stage, which trains LLMs on human-written or GPT-synthesized (instruction, response) pairs [40, 29]. This paradigm neglects the discrepancy between pretraining and SFT data, potentially inducing hallucinations by instructing LLMs to appear helpful, even when they are unable to solve the problem, and discouraging them from expressing uncertainty or declining responses [47]. In this work, we propose SaySelf to train LLMs to express accurate confidence estimates accompanied with their responses.
LLMs’ Confidence Elicitation
Eliciting accurate confidence estimates for LLM-generated answers that contain multiple tokens is challenging [4, 50]. Previous work can be categorized into prompting-based and training-based approaches. Prompting-based approaches use a specific prompt to guide LLMs to generate confidence scores for their predictions [36, 18], or prompt LLMs to generate the answers multiple times and use the consistency levels as indicators of their confidence [42, 26]. These approaches can cause inferior performance or lead to extensive extra inference-time latency. Training-based approaches construct a specialized dataset for supervised fine-tuning, encouraging LLMs to express their uncertainty. [22] first group examples based on their types of question as the label, then obtains the confidence score for each example using the empirical accuracy for the whole group. This approach can lead to suboptimal confidence estimates since not all examples in the same group are equal. R-Tuning [47] reconstructs the SFT data to add “I am sure/unsure” at the end of the correct/incorrect responses, which can only generate binary uncertainty estimates. We show that SaySelf guides LLMs to generate more accurate and fine-grained confidence estimates.
@@
Our work is also related to explainability for LLMs regarding the self-reflective rationales generation [49, 33]. Previous work on natural language explanations for LLMs provides motivation to explain the models’ decision-making process for a prediction [8, 5]. The typical approaches to producing natural language explanations involve training LLMs with the ground-truth labels and the human-annotated explanations that can serve as effective augmented supervision that guide LLMs to reason in a right way [30, 24, 46]. SaySelf significantly departs from existing methods by generating rationales that not only justify the predictions but also elucidate the confidence estimates. Most importantly, SaySelf adopts LLMs’ internal reasoning process to generate self-reflective rationales, instead of human-annotated explanations, which may not be faithful to specific LLMs. Another line of research adopts chain-of-thought (CoT) reasoning as natural language explanations [41, 25, 43, 6]. Unlike CoT, which primarily clarifies the rationale behind predictions, our method also explicates the sources of uncertainty.
@@ 3 SaySelf
We present SaySelf, a training framework to teach LLMs to express fine-grained confidence with self-reflective rationales (see Figure 2). SaySelf consists of 2 essential stages: (1) Supervised Fine-Tuning: We establish a model-specific dataset containing self-reflective rationales and confidence estimates. This dataset is constructed based on multiple sampled responses from LLMs. (2) Reinforcement Learning from Task Supervision: We use reinforcement learning with a carefully designed reward function to further calibrate the confidence estimates for each instance. For both 2 stages, we adopt the training samples in HotpotQA [44], which typically require multi-step reasoning on knowledge facts to derive the answer.
@@ 3.1 Supervised Fine-Tuning
In this stage, our goal is to construct a supervised dataset $D$, where each sample contains a question $q$, an answer with the reasoning chain $s$, the self-reflective rationale $r$, and the confidence estimate $c$. Basically, $r$ is a summary of uncertainty in specific knowledge, and is generated by analyzing the inconsistency in multiple selective responses sampled from the vanilla LLM $M$. $c$ is an integer from 1 to 10, and is derived based on the consistency of $s$.
We adopt 90K questions in HotpotQA. For each question, we prompt $M$ to generate the reasoning chain and the answer $N$ times. We perform clustering on the $N$ responses to obtain $K$ representative clusters based on the semantic similarity among responses since there is significant redundancy. Specifically, we adopt the Instructor [34], an instruction-finetuned text embedding model that produces text embeddings customized to the specific task and domain. Our clustering process involves examining each response, identifying those within the similarity threshold $T$, and grouping them accordingly until all responses have been processed. The cluster size $S$ is defined as the number of responses in the cluster. We randomly pick one selected response per cluster for the following steps, as empirical evidence suggests significant similarity among responses within the same cluster.
To derive the confidence estimate $c$, we first check the correctness of the selected response from each group using the golden answer annotated in HotpotQA. Samples with none correct response are removed since we should avoid training LLMs on incorrect examples. The correct response is taken as the golden $s$ for this sample, and $c$ is computed as: $c=round(\frac{S_{c}}{N}*10)$, where $S_{c}$ is the size of $s$’s cluster, and $round(x)$ returns the nearest integer of $x$
@@
$\max_{\Theta}\sum_{(q,s,r,c^{\prime})\in D}\log P(s|q;\Theta)+\log P(r|s,q;\Theta)+\log P(c^{\prime}|s,r,q;\Theta)$
where $\Theta$ represents the parameters $M$, $c^{\prime}$ is the natural language expression of the confidence estimate $c$ (a.k.a., My confidence is $c$).
@@ 3.2 Reinforcement Learning from Task Supervision
Due to the nature of supervised fine-tuning, the model tends to produce homogeneous confidence levels, such as relatively lower confidence levels for correct responses and higher levels for incorrect responses. To address this issue, we use reinforcement learning to further calibrate LLMs’ fine-grained confidence estimates and encourage the model to produce more accurate and differentiated values. During the sampling phase, LLMs are prompted to produce responses, self-reflective rationales, and confidence levels. To optimize, we assess the generated response against the ground truth. Subsequently, we formulate a reward function considering answer accuracy and model confidence. To encourage the model towards more differentiated values, the reward function has a quadratic output:
$R=\frac{((I(\text{correct})\times\text{confidence level})-(I(\text{incorrect})\times\text{confidence level}))^{2}}{100}$
where $I()$ is the indicator function, return 1 if the condition is True, else 0. Basically, this function is designed to reinforce LLMs for high confidence in accurate samples while penalizing them for overconfident in incorrect ones.
We utilize the Proximal Policy Optimization (PPO) algorithm [32] to train LLMs based on this defined reward function. The optimization objective is expressed as:
$\max_{\Theta}\ \ \mathbb{E}_{t}\left[\min(r_{t}(\Theta)\hat{A}_{t},\text{clip}(r_{t}(\Theta),1-\epsilon,1+\epsilon)\hat{A}_{t})\right]$
where $r_{t}(\Theta)$ calculates the probability ratio of the newly proposed policy to the old policy. The advantage estimate $\hat{A}_{t}$, crucial for directing updates, is calculated from the difference between the anticipated future rewards under the current policy and the baseline or value function. This advantage is directly influenced by the reward $R$, which in turn ties the optimization process closely with both response accuracy and confidence level.
@@ 3.3 Implementation Details
For the supervised fine-tuning, the sampling time $N$ is set to 100 and the temperature is 1.2. The similarity threshold $T$ is set to 0.9. For the reinforcement learning stage, the learning rate $\mu$ is set to 1e-5 and the batch size $B$ is set to 8. To check the correctness of the responses, we verify that the annotated answers are included within them. This heuristic approach has been empirically shown to yield high precision in knowledge-based QA tasks.
@@ 4.1 Evaluation Setting
Following R-Tuning [47], we evaluate LLMs on the knowledge-extensive QA tasks. We include the following datasets for evaluation: HotpotQA [44], a dataset of multi-hop reasoning question-answer pairs; TruthfulQA [21], a dataset that tests whether models generate truthful answers to questions specifically designed to induce false answers; StrategyQA [11], a dataset of true/false questions requiring multi-hop reasoning; FEVER [35], a dataset used to assess the ability of models to verify the factuality of statements against Wikipedia documents; HaluEval [19], a dataset that evaluates the hallucination rate of models; ParaRel [10], a dataset that measures the model’s performance in understanding paraphrased relational facts.
@@
We measure various approaches from 3 aspects. (1) Confidence Calibration Performance: We adopt 2 calibration metrics. First, we use the ECE score to measure the confidence calibration error [12, 7]. Basically, ECE evaluates the correlation between the confidence scores assigned by LLMs and their corresponding correctness. Second, we adopt the AUROC score following [14]. It measures the ability of LLMs to distinguish between correct and incorrect responses across different threshold settings. (2) Task Performance: We measure the typical accuracy on the test split of the datasets. (3) Faithfulness of the Generated Self-Reflective Rationales: We make the first effort to measure the faithfulness of the provided self-reflective rationales. We suggest employing the same intuition utilized in SaySelf. For each question, we sample multiple responses (answers with reasoning chains) from the LLM, and perform clustering to retain several representative responses. Subsequently, we utilize a proficient LLM (GPT-4) to examine whether the provided self-reflective rationales can faithfully express the uncertainty demonstrated in the sampled responses, and give a score from 1 to 10. The final faithfulness score is the average over all rationales.
Table 1: The ECE evaluation results of baselines, SaySelf, and various ablations. Lower is better. HotpotQA is the only in-distribution dataset.
@@
w/o R 0.3833 0.4308 0.4125 0.3973 0.4344 0.3926
Reward function 0.6129 0.4356 0.4062 0.4238 0.2812 0.3316
Table 2: The AUROC evaluation results of baselines, SaySelf, and various ablations. HotpotQA is the only in-distribution dataset.
DP 0.3222 0.5667 0.5193 0.5371 0.5278 0.5291
SC 0.5765 0.4939 0.5498 0.5472 0.5843 0.5546
R-Tuning 0.6529 0.5980 0.5406 0.5688 0.5330 0.5424
GCE 0.5042 0.4966 0.5043 0.4942 0.4907 0.5031
SaySelf 0.7156 0.6107 0.6074 0.6481 0.7318 0.6816
w/o RL 0.6524 0.5675 0.5910 0.5798 0.5929 0.6003
w/o R & CE 0.5256 0.5724 0.5738 0.6059 0.6002 0.5823
w/o R 0.4955 0.4835 0.5391 0.5101 0.5569 0.5713
Reward function 0.5140 0.4907 0.5091 0.5137 0.5147 0.5053
Table 3: The accuracy evaluation results of baselines, SaySelf, and various ablations. HotpotQA is the only in-distribution dataset.
@@
w/o R 0.3379 0.4861 0.3670 0.4539 0.7547 0.5684
Reward function 0.3715 0.5721 0.5811 0.5443 0.8450 0.6577
Table 4: The faithfulness evaluation results for self-reflective rationales. HotpotQA is the only in-distribution dataset.
@@
We compare with the following approaches: (1) Direct prompting for confidence extraction (DP): We directly ask the vanilla LLM to give a confidence score from 1 to 10 in their previous response [36]. (2) Self-consistency-based confidence estimate (SC): We use the self-consistency-based approach to derive the confidence estimates of LLMs. Confidence is calculated as the ratio of response frequency to the sampling time [42]. (3) R-Tuning: We directly train LLMs to generate binary confidence estimates (sure vs. unsure) using a model-specific dataset [47]. (4) Grouping-based confidence estimates for calibration training (GCE): We group the samples in HotpotQA via clustering, and use the accuracy of samples in the group as the confidence estimates for all samples within that group. The constructed dataset is thus used for fine-tuning [22]. We implement the baseline approaches and SaySelf on the Mistral LLM [17] for fair comparison.

### v3

@@ Abstract
Large language models (LLMs) often generate inaccurate or fabricated information and generally fail to indicate their confidence, which limits their broader applications. Previous work has elicited confidence from LLMs by direct or self-consistency prompting, or constructing specific datasets for supervised finetuning. The prompting-based approaches have inferior performance, and the training-based approaches are limited to binary or inaccurate group-level confidence estimates. In this work, we present SaySelf, a novel training framework that teaches LLMs to express more fine-grained confidence estimates. In addition, beyond the confidence scores, SaySelf initiates the process of directing LLMs to produce self-reflective rationales that clearly identify gaps in their parametric knowledge and explain their uncertainty. This is achieved by using an LLM to automatically summarize the uncertainties in specific knowledge via natural language. The summarization is based on the analysis of the inconsistency in multiple sampled reasoning chains, and the resulting data is utilized for supervised fine-tuning. Moreover, we utilize reinforcement learning with a meticulously crafted reward function to calibrate the confidence estimates, motivating LLMs to deliver accurate, high-confidence predictions and to penalize overconfidence in erroneous outputs. Experimental results demonstrate the effectiveness of SaySelf in reducing the confidence calibration error and maintaining the task performance. The generated self-reflective rationales are also reasonable and can further contribute to the calibration. The code is made public at https://github.com/xu1868/SaySelf.
SaySelf: Teaching LLMs to Express Confidence
with Self-Reflective Rationales
Tianyang Xu1thanks: Equal contribution., Shujin Wu3∗, Shizhe Diao4, Xiaoze Liu1 Xingyao Wang2, Yangyi Chen2thanks: Corresponding author., Jing Gao1† 1 Purdue University 2 University of Illinois Urbana-Champaign 3 University of Southern California 4 NVIDIA {xu1868, xiaoze, jinggao}@purdue.edu, shujinwu@usc.edu sdiao@nvidia.com, {xingyao6,yangyic3}@illinois.edu
@@ 1 Introduction
While large language models (LLMs) exhibit remarkable proficiency in reasoning and generating effective responses OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Wang et al. (2024), they often produce fabricated information (a.k.a, hallucination) and typically hesitate to indicate their uncertainty when faced with unfamiliar questions Ye et al. (2023); Liu et al. (2023). Determining how to accurately obtain reliable confidence estimates from LLMs is essential Xiong et al. (2023); Zhou et al. (2023), particularly when the responses are not limited to single tokens [^1] The confidence of a single-token answer can be derived from its token probability, which typically exhibits high calibration OpenAI (2023); Chen et al. (2023b)..
@@
Previous work on eliciting confidence from LLMs includes prompting-based and training-based approaches. Prompting-based methods, such as direct prompting and self-consistency prompting in Figure 1, employ specific prompts to generate confidence scores or use answer consistency as a confidence indicator, even though these can have poor calibration performance or significantly increase inference latency Tian et al. (2023); Xiong et al. (2023); Diao et al. (2023). Training-based approaches, such as group-based calibration training and R-Tuning in Figure 1, develop specialized datasets for fine-tuning that encourage LLMs to express confidence. However, these methods often provide suboptimal or binary confidence estimates, failing to accurately reflect the models’ confidence levels Lin et al. (2022); Zhang et al. (2023a); Yang et al. (2023). In conclusion, previous work tends to suffer from the following problems: (1) Poor calibration performance; (2) Coarse-grained confidence levels; (3) Long inference latencies.
@@
In this work, we present SaySelf, a training framework that teaches LLMs to generate more accurate and fine-grained confidence estimates. It successfully tackles the aforementioned problems in previous work. More importantly, SaySelf also goes beyond the confidence elicitation in previous work, and further enables LLMs to generate self-reflective rationales that indicate their knowledge gap and explain their confidence estimates (Figure 1). We accomplish this by automatically generating a model-specific dataset for supervised fine-tuning using an off-the-shelf LLM (e.g., GPT-4 OpenAI (2023)). Specifically, for each question, we sample multiple reasoning chains from LLMs. We then perform clustering of the reasoning chains based on the semantic similarity and retain one instance per cluster. GPT-4 is then tasked with analyzing these instances from various clusters, summarizing uncertainties in natural language from a first-person perspective, which is subsequently used for fine-tuning.
@@
For accurate and fine-grained confidence estimates, we employ reinforcement learning to calibrate LLMs’ confidence estimate in each response. We design a reward function that incentivizes LLMs to produce accurate, high-confidence predictions and imposes penalties for overconfidence in incorrect responses. In addition, the self-reflective rationales and confidence estimates are generated without multiple sampling, significantly reducing the inference time.
@@
We evaluate SaySelf on multiple knowledge-extensive question-answering tasks. We show that SaySelf significantly reduces the confidence calibration error and maintains the task performance. The generated self-reflective rationales effectively capture the internal uncertainty and can further improve the calibration performance. In addition, we find that SaySelf enables LLMs to produce low confidence in unanswerable questions.
@@ 2 Related Work
LLMs’ Confidence Elicitation
Eliciting accurate confidence estimates for LLM-generated answers that contain multiple tokens is challenging Borji (2023); Zhou et al. (2024). Previous work can be categorized into prompting-based and training-based approaches. Prompting-based approaches use a specific prompt to guide LLMs to generate confidence scores for their predictions Tian et al. (2023); Kadavath et al. (2022), or prompt LLMs to generate the answers multiple times and use the consistency levels as indicators of their confidence Xiong et al. (2023); Lyu et al. (2024); Diao et al. (2023); Yang et al. (2023). These approaches can cause inferior performance or lead to extensive inference-time latency. Training-based approaches construct a specialized dataset for supervised fine-tuning, encouraging LLMs to express their uncertainty. Lin et al. (2022) first group examples based on their types of question as the label, then obtains the confidence score for each example using the empirical accuracy for the whole group. This approach can lead to suboptimal confidence estimates since not all examples in the same group are equal. R-Tuning Zhang et al. (2023a) reconstructs the SFT data to add “I am sure/unsure” at the end of the correct/incorrect responses, which can only generate binary uncertainty estimates. As mentioned in Section 1, SaySelf addresses the limitations of the previous methods and guides LLMs to generate more accurate and fine-grained confidence estimates.
@@
LLMs’ hallucination refers to instances where these models generate information that is not supported by their training data or the input provided Zhang et al. (2023b); Liang et al. (2024); Agrawal et al. (2023). Numerous research is dedicated to exploring the causes of hallucination Dziri et al. (2022); McKenna et al. (2023); Han et al. (2024) and developing methods to detect or mitigate hallucination Varshney et al. (2023); Rawte et al. (2023); Andriopoulos and Pouwelse (2023). Besides the hallucination, the reluctance of LLMs to express uncertainty when they are unable to solve tasks can further erode trust in these systems Ji et al. (2023); Zhou et al. (2024). Existing research identifies the tendency in LLMs to fabricate information when addressing unknown questions Liu et al. (2023); Hu et al. (2023); Amayuelas et al. (2023). This inability can be traced back to the supervised instruction finetuning (SFT) stage, which trains LLMs on human-written or GPT-synthesized (instruction, response) pairs Wang et al. (2022); Peng et al. (2023). This paradigm neglects the discrepancy between pretraining and SFT data, potentially inducing hallucinations by instructing LLMs to appear helpful, even when they are unable to solve the problem, and discouraging them from expressing uncertainty or declining responses Zhang et al. (2023a). In this work, we propose SaySelf to train LLMs to express accurate confidence estimates and self-reflective rationales as an efficient method against hallucinations, as they can guide end users to verify information in responses and help them regain confidence in LLMs.
@@
Our work is also related to explainability for LLMs regarding the self-reflective rationales generation Zhao et al. (2024); Singh et al. (2024). Previous work on natural language explanations for LLMs provides motivation to explain the models’ decision-making process for a prediction Costa et al. (2018); Cambria et al. (2023). The typical approaches to producing natural language explanations involve training LLMs with the ground-truth labels and the human-annotated explanations that can serve as effective augmented supervision that guide LLMs to reason in a right way Rajani et al. (2019); Luo et al. (2021); Yordanov et al. (2021). Another line of research adopts chain-of-thought (CoT) reasoning as natural language explanations Wei et al. (2022); Lyu et al. (2023); Xu et al. (2024); Chen et al. (2023a). Compared to previous methods, SaySelf significantly departs from existing methods by generating rationales that not only justify the predictions but also elucidate the confidence estimates. Most importantly, SaySelf adopts LLMs’ internal reasoning process to generate self-reflective rationales, instead of human-annotated explanations, which may not be faithful to specific LLMs. Unlike CoT, which primarily clarifies the rationale behind predictions, SaySelf also explicates the sources of uncertainty.
@@ 3 SaySelf
We present SaySelf, a training framework to teach LLMs to express fine-grained confidence with self-reflective rationales (see Figure 2). SaySelf consists of 2 essential stages: (1) Supervised Fine-Tuning: We establish a model-specific dataset containing self-reflective rationales and confidence estimates. This dataset is built from multiple sampled responses from LLMs. (2) Reinforcement Learning from Task Supervision: We use reinforcement learning with a carefully designed reward function to further calibrate the confidence estimates for each instance. For both 2 stages, we adopt the training samples in HotpotQA Yang et al. (2018), which typically require multi-step reasoning on knowledge facts to derive the answer. After the two-stage training, the trained models can directly answer questions with confidence estimates and self-reflective rationales without additional computational overhead.
@@ 3.1 Supervised Fine-Tuning
In this stage, our goal is to construct a supervised dataset $D$, where each sample contains a question $q$, an answer with the reasoning chain $s$, the self-reflective rationale $r$, and the confidence estimate $c$. Basically, $r$ summarizes specific knowledge that the LLM is uncertain about, and is generated by analyzing the inconsistency in multiple selective responses sampled from the vanilla LLM $M$. $c$ is an integer from 1 to 10, and is derived based on the consistency of $s$.
We adopt 90K questions in HotpotQA. For each question, we prompt $M$ to generate the reasoning chain and the answer $N$ times. We perform clustering on the $N$ responses to obtain $K$ representative clusters based on the semantic similarity among responses since there is significant redundancy. Specifically, we adopt the Instructor Su et al. (2022), an instruction-finetuned text embedding model that produces text embeddings customized to the specific task and domain. Our clustering process involves examining each response, identifying those within the similarity threshold $T$, and grouping them accordingly until all responses have been processed. The cluster size $S$ is defined as the number of responses in the cluster. We randomly pick one selected response per cluster for the following steps, as empirical evidence suggests significant similarity among responses within the same cluster (see Appendix B for details).
To derive the confidence estimate $c$, we first check the correctness of the selected response from each group using the golden answer annotated in HotpotQA. Samples with no correct responses are removed to avoid training LLMs on incorrect examples. The correct response is taken as the golden $s$ for this sample, and $c$ is computed as: $c=round(\frac{S_{c}}{N}*10)$, where $S_{c}$ is the size of $s$’s cluster, and $round(x)$ returns the nearest integer of $x$.
@@
$\max_{\Theta}\sum_{(q,s,r,c^{\prime})\in D}[\ \log P(s|q;\Theta)$
$+\log P(r|s,q;\Theta)+\log P(c^{\prime}|s,r,q;\Theta)]$ (1)
where $\Theta$ represents the parameters of $M$, $c^{\prime}$ is the natural language expression of the confidence estimate $c$ (a.k.a., “My confidence is $c$”). The objective function is meant to maximize the sum of these log probabilities over all the tuples $(q,s,r,c^{\prime})$ in the dataset $D$.
@@ 3.2 Reinforcement Learning from Task Supervision
Due to the nature of supervised fine-tuning, the model tends to produce homogeneous confidence levels, such as relatively lower confidence levels for correct responses and higher levels for incorrect responses. To address this issue, we use reinforcement learning to further calibrate LLMs’ fine-grained confidence estimates and guide the model to produce more accurate and differentiated values. During the sampling phase, LLMs are prompted to produce responses, self-reflective rationales, and confidence levels. To optimize the model, we compare the generated response with the ground truth. Subsequently, we formulate a reward function considering answer accuracy and model confidence. To encourage the model towards more differentiated values, the reward function has a quadratic output:
$R=1-2*(\mathbb{I}(\textrm{response})-\textrm{confidence level})^{2}$ (2)
where $\mathbb{I}()$ is the indicator function, which returns 1 if the generated response is correct, else 0. The confidence level is normalized between 0 and 1. This reward function reinforces LLMs for high confidence in accurate samples while penalizing them for being overconfident in incorrect ones.
We utilize the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017) to train LLMs based on this defined reward function. The optimization objective is expressed as:
$\max_{\Theta}\ \ \mathbb{E}_{t}[$ $\min(r_{t}(\Theta)\hat{A}_{t},\text{clip}(r_{t}(\Theta),$
$1-\epsilon,1+\epsilon)\hat{A}_{t})]$ (3)
where $r_{t}(\Theta)$ calculates the probability ratio of the newly proposed policy to the old policy. The advantage estimate $\hat{A}_{t}$, crucial for directing updates, is calculated from the difference between the anticipated future rewards under the current policy and the baseline or value function. This advantage estimate is directly influenced by the reward $R$, which in turn ties the optimization process closely with both response accuracy and confidence level.
@@ 3.3 Implementation Details
For the supervised dataset collection, the sampling time $N$ is set to 100 and the temperature is 1.2. The similarity threshold $T$ is set to 0.9. For supervised fine-tuning, the learning rate is set to 7e-5 and the batch size is set to 8. For the reinforcement learning stage, the learning rate is set to 1e-5 and the batch size is set to 8. To check the correctness of the responses, we utilize a verification method where annotated answers must be present within the responses. This heuristic demonstrates high precision in knowledge-based QA tasks.
@@ 4.1 Evaluation Setting
We follow Zhang et al. (2023a) to evaluate LLMs on knowledge-extensive QA tasks. We include the following datasets: HotpotQA Yang et al. (2018), a dataset of multi-hop reasoning question-answer pairs; TruthfulQA Lin et al. (2021), a dataset that tests whether models generate truthful answers to questions specifically designed to induce false answers; StrategyQA Geva et al. (2021), a dataset of true/false questions requiring multi-hop reasoning; FEVER Thorne et al. (2018), a dataset used to assess the ability of models to verify the factuality of statements against Wikipedia documents; HaluEval Li et al. (2023), a dataset that evaluates the hallucination of models; ParaRel Elazar et al. (2021), a dataset that measures the model’s performance in understanding paraphrased relational facts.
Evaluation Environments
The experiments are run on a server with 4 Nvidia A6000 GPUs and 256GB RAM. The models are implemented with the Huggingface Transformers (https://huggingface.co/) library. The reported data are all average values of three runs. Both stages take approximately 1 hour to train during the two-stage training process.
@@
We measure various approaches from 3 aspects. (1) Confidence Calibration Performance: We adopt 2 calibration metrics. First, we use the ECE score to measure the confidence calibration error Guo et al. (2017); Chen et al. (2023b). Basically, ECE evaluates the correlation between the confidence scores assigned by LLMs and their corresponding correctness. For responses from LLMs $A$, it can be calculated as
$\mathrm{ECE}=\frac{1}{|A|}\sum_{a\in A}|\mathbb{I}(a)-\mathrm{conf}(a)|,$ (4)
where $\mathbb{I}()$ is the indicator function defined in Equation 2, and $\mathrm{conf}()$ returns the confidence level of LLMs. Second, we adopt the AUROC score following Hendrycks and Gimpel (2016). It measures the ability of LLMs to distinguish between correct and incorrect responses across different threshold settings. It can be calculated as
$\mathrm{AUROC}=\int_{0}^{1}\mathrm{TPR}(\mathrm{FPR^{-1}}(x))\mathrm{d}x,$ (5)
where $x$ is the threshold confidence level, $\mathrm{TPR}$ is the true positive rate under this threshold confidence level, and $\mathrm{FPR}$ is the false positive rate under the threshold. (2) Task Performance: We measure the typical accuracy on the test split of the datasets. (3) Faithfulness of the Generated Self-Reflective Rationales: We make the first effort to measure the faithfulness of the provided self-reflective rationales. We suggest employing the same intuition utilized in SaySelf. For each question, we sample multiple responses (answers with reasoning chains) from the LLM, and perform clustering to retain several representative responses. Subsequently, we utilize a proficient LLM (GPT-4) to examine whether the provided self-reflective rationales can faithfully express the uncertainty demonstrated in the sampled responses, and give a score from 1 to 10. The final faithfulness score is the average over all samples.
@@
w/o R 0.3750 0.3609 0.3938 0.3854 0.4294 0.4730
w/ Naive RF 0.6129 0.4356 0.4062 0.4238 0.2812 0.3316
Table 1: The ECE evaluation results of baselines, SaySelf, and various ablations. Lower is better. HotpotQA is the only in-distribution dataset. $p$-Values are the $p$-values comparing SaySelf over other methods. In this table, DP denotes direction prompting, SC denotes self-consistency, PC denotes prompting for correctness, AS denotes aligning with self-consistency-based confidence, GCE denotes grouping-based confidence estimates for calibration training; w/o RL denotes SaySelf without reinforcement learning, w/o R & CE denotes SaySelf without self-reflective rationales and confidence estimates, w/o R denotes SaySelf without self-reflective rationales, w/ Naive RF denotes using another naive reward function. The numbers with asterisk marks (*) mean significant advantage with the statistical significance threshold of $p$-value 0.05 in the paired t-test comparing with baselines.
@@
w/o R 0.3650 0.4964 0.4224 0.4848 0.7652 0.5707
w/ Naive RF 0.3715 0.5721 0.5811 0.5443 0.8450 0.6577
Table 2: The accuracy evaluation results of baselines, SaySelf, and various ablations.
@@
Table 3: The faithfulness evaluation results for self-reflective rationales.
@@
We compare with the following approaches: (1) Direct prompting for confidence extraction (DP): We directly ask the vanilla LLMs to give a confidence score from 1 to 10 in their previous response Tian et al. (2023). (2) Self-consistency-based confidence estimate (SC): We use the self-consistency-based approach to derive the confidence estimates of LLMs. Confidence is calculated as the ratio of response frequency to the number of samples Xiong et al. (2023). (3) Prompting for correctness (PC): We ask the vanilla LLMs to judge whether their responses are correct or not (Kadavath et al., 2022). (4) R-Tuning: We train LLMs to generate binary confidence estimates (sure vs. unsure) using a model-specific dataset Zhang et al. (2023a). (5) Aligning with self-consistency-based confidence (AS): We train LLMs to generate the confidence estimates derived from self-consistency prompting (Yang et al., 2023). (6) Grouping-based confidence estimates for calibration training (GCE): We group the samples in HotpotQA via clustering, and use the accuracy of samples in the group as the confidence estimates for all samples within that group. The constructed dataset is thus used for fine-tuning Lin et al. (2022). We implement the baseline approaches and SaySelf on Mistral-7B Jiang et al. (2023) for fair comparison. To prove that SaySelf can generalize on multiple models, we also implement the baseline approaches and SaySelf on Llama 3 8B Team (2024) in Appendix D.